OpenStudy (anonymous):

which equatiom represents y=-x^2-10x-20 in vertex form a.y=-(x-5)^2+15 b.y=-(x+5^2+5 c.y=-(x+5)^2+10 d.y=-(x-5)^2+5

OpenStudy (anonymous):

y=-x^2-10x-20 y=-(x^2+10x+25)+25-20 y=-(x+5)^2+5 so u get b. understood ?

OpenStudy (anonymous):

yeah